The width of a rectangle is 3 inches more than one-half of its length. If the perimeter is 60 inches, find the length and width of the rectangle.
I'm unsure of how to get the formula of how to do this.

Answer by addingup(3677)

the width is...
W =
3 inches more than one-half its length
W = 1/2L+3
The perimeter is 2L+2W and the problem tells you that it's 60
So far, notice that all the information is given by the problem.
-------------------------------------------------------
2L+2W = 60 substitute for W, with the value given by the problem:
2L+2(1/2L+3) = 60
Distribute the 2 (Distributive Properties of Multiplication)
2L+L+6 = 60
3L= 54
L = 54/3 = 18
now that we know the length, find the width. The problem says:
W = 1/2L+3
W = 1/2(18)+3 What's one-half of 18? Use it next:
W = 9+3 = 12
Check:
2L + 2W = 60
2(18) + 2(12) = 60
36 + 24 = 60 add the left side and you'll see that I have the correct answer.
:
Happy learning,
John